OpenStudy (anonymous):

]Choose the equation of the line passing through the point (4, -2) and perpendicular to y = 1/3x - 7. y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

OpenStudy (anonymous):

I think there is a property regarding the slopes of perpendicular lines: the product of the slopes = -1. So, just going off of that we can eliminate two of your options right off the bat. Try to take it from there and see if you remember the rest.

OpenStudy (anonymous):

all of the slopes are 3 so i dont know what your trying to say

OpenStudy (anonymous):

y=mx+b M=slope

OpenStudy (anonymous):

If the product of the slopes must equal -1, and the slope of your original equation is positive, then what does that mean regarding the sign of the slopes you have to choose from?

OpenStudy (anonymous):

okay got you now but i still dont know what the answer is nor how to do it

OpenStudy (anonymous):

There may be other ways to do this, but off the top of my head (without using a graphing calculator), I am thinking that if the lines are perpendicular then that means they only intersect at one point, right? So, couldn't we just set our original equation to equal one of the equations with a negative slope? If they meet at (4, -2) then that is the right choice.

OpenStudy (anonymous):

flutter it im guessing

OpenStudy (anonymous):

* f uc k

OpenStudy (anonymous):

don't. try this: plug in the coordinates (4, -2) to the two functions we think it might be. For example, with the first one y = -3x -14, it would be: -2 = -3(4) -14 then evaluate it to see if it is true.

OpenStudy (anonymous):

I'll spare you the suspense, the one I just wrote ISN'T the right one. Do what I did, but with y = -3x + 10